OpenStudy (anonymous):

The Cafeteria has a fixed price for the salad bar per ounce. Sandwich cost $3.10. Drink cost $1.75. If a 7-ounce and a drink that cost the same as a 4-ounce salad and a sandwich, how much does the salad cost per ounce?

OpenStudy (anonymous):

The salad costs .45 per ounce.

OpenStudy (anonymous):

s=cost of salad per ounce 7s + 1.75 = 4s + 3.10 -4s -4s _________________________ 3s +1.75 = 3.10 -1.75 -1.75 ___________________ 3s = 1.35 s = .45